# Improving the Distributional Alignment of LLMs using Supervision

### v2

@@ A.1 Alignment results for all models
A.2 Log Probability Results
We get the model’s log probabilities for each answer choice Geng et al. (2024) and normalize to get the distribution over all answer choices, following prior work Santurkar et al. (2023). We obtain log probabilities for the smaller Llama models. Log probability results are shown in Table 6.
@@ A.2 Log Probability Results
Table 7: Opinion alignment before and after calibration for each dataset, LLM, and elicitation method. Each pair of columns compares the base-generated or SD-generated distributions to the calibrated distributions ($C$) for each elicitation method: paraphrase (‘P’), self-random (‘S’), and verbalized (‘V’). Bolded values are significant between each pair. The mean and standard deviation across models are shown in the bottom rows of each dataset section. We note that for some verbalized elicitation settings, Llama-3.2-1B was unable to follow instructions and either did not return any distributions in the format we required for parsing, or refused to produce a distribution. See Table 17.
A.3 Individual Sociodemographic Results
@@ A.4 Out-of-domain generalization
Table 11: Opinion alignment before and after calibration for each dataset, LLM, and elicitation method, when training on two datasets and evaluating on the out-of-domain dataset. Each pair of columns compares the base-generated or SD-generated distributions to the calibrated distributions ($C$) for each elicitation method: paraphrase ($P$), self-random ($S$), and verbalized ($V$). Bolded values are significant between each pair. The mean and standard deviation across models are shown in the bottom rows of each dataset section.
@@
Table 12: Opinion alignment before and after calibration for each dataset and LLM using log probability distributions, when training on two datasets and evaluating on the out-of-domain dataset. Each pair of columns compares the base-generated or SD-generated distributions to the calibrated distributions ($C$) for log probabilities ($L$). Bolded values are significant between each pair. The mean and standard deviation across models are shown in the bottom rows of each dataset section.
Appendix B Prompts (Reproducibility)
B.1 Demographics used
Demographics are shown in Table 13.
Dataset Qs Example question Demographics
How much do you trust doctors and nurses in your country?
B.2 Base prompts
Base prompts are shown in Table 14.
Prompt type Prompt template
B.3 Sociodemographic prompts
Sociodemographic prompts are shown in Table 15.
Prompt type Prompt template
B.4 Prompt variations
Prompt variations are shown in Table 16.
Prompt variation Prompt template
Table 16: Prompt templates for the different prompt variants. This text is combined with the templates for both base prompts (Table 14) and SD prompts (Table 15) to get the results shown in Tables 2 and 4. All other results use the standard prompts without variation.
Appendix C Extracted distributions

### v3

@@ D.1 Alignment results for all models
D.2 Log Probability Results
We get the model’s log probabilities for each answer choice Geng et al. (2024) and normalize to get the distribution over all answer choices, following prior work Santurkar et al. (2023). We obtain log probabilities for the smaller Llama models. Log probability results are shown in Table 6. We find that using log probabilities with calibration yielded the most aligned distributions in the 10/24 settings considered. Future work might look more into comparisons between logit-based and verbalized distribution elicitation.
@@ D.2 Log Probability Results
Table 7: Opinion alignment before and after calibration for each dataset, LLM, and elicitation method. Each pair of columns compares the base-generated or SD-generated distributions to the calibrated distributions ($C$) for each elicitation method: paraphrase (‘P’), self-random (‘S’), and verbalized (‘V’). Bolded values are significant between each pair (calculated via paired t-test and Bonferroni correction, see Appendix G). The mean and standard deviation across models are shown in the bottom rows of each dataset section. We note that for some verbalized elicitation settings, Llama-3.2-1B was unable to follow instructions and either did not return any distributions in the format we required for parsing, or refused to produce a distribution (see Table 13).
D.3 Individual Sociodemographic Results
@@ D.4 Out-of-domain generalization
Table 11: Opinion alignment before and after calibration for each dataset, LLM, and elicitation method, when training on two datasets and evaluating on the out-of-domain dataset. Each pair of columns compares the base-generated or SD-generated distributions to the calibrated distributions ($C$) for each elicitation method: paraphrase (‘P’), self-random (‘S’), and verbalized (‘V’). Bolded values are significant between each pair. The mean and standard deviation across models are shown in the bottom rows of each dataset section.
@@
Table 12: Opinion alignment before and after calibration for each dataset and LLM using log probability distributions, when training on two datasets and evaluating on the out-of-domain dataset. Each pair of columns compares the base-generated or SD-generated distributions to the calibrated distributions ($C$) for log probabilities (‘L’). Bolded values are significant between each pair. The mean and standard deviation across models are shown in the bottom rows of each dataset section.
Appendix E Extracted distributions
@@
Appendix F Minimal supervision training plots
We plot Mean Squared Error (MSE) over training data size in Figure 4, showing that MSE usually converges between 1 and 10 examples, though this is model and dataset dependent.
Appendix G Statistical significance details
We list the t-test statistic and p-value after Bonferroni correction in Table 14 for our main results (Tables 1 and 7). Our Bonferroni alpha is 0.000172, calculated using the Python Statsmodels [^4] https://www.statsmodels.org/stable/index.html package. Our results are calculated on the 10% test split discussed in Section 3.6. While the number of examples is different for each of the 270 settings, the average number of examples is approximately 237. While we do not conduct explicit tests for normality for using the t-test, the t-test is remarkably robust to violations in sample normality Smucker et al. (2007), particularly given the number of examples.
Model t-test statistic Corrected p-value
$P_{B}$ $P_{S}$ $_{B} SS$$_{S}$ $V_{B}$ $V_{S}$ $P_{B}$ $P_{S}$ $_{B} SS$$_{S}$ $V_{B}$ $V_{S}$
WGM OLMo-2-7B-Base 11.76 5.75 3.07 7.76 1.21 4.99 2.70e-24 6.04e-06 6.74e-01 3.33e-11 1.00 3.59e-04
OLMo-2-7B-SFT 1.18 6.01 11.11 9.34 15.54 13.24 1.00 1.50e-06 5.57e-22 5.00e-16 1.91e-38 1.36e-29
OLMo-2-7B-DPO 4.20 8.99 18.04 10.32 16.17 14.32 1.01e-02 6.41e-15 3.94e-48 2.84e-19 6.95e-41 2.33e-33
OLMo-2-7B-Instruct 16.07 12.93 13.79 13.32 17.65 9.12 1.79e-40 1.67e-28 9.17e-32 5.70e-30 1.14e-42 4.05e-15
Llama-3-70B 20.48 19.53 8.47 22.37 5.37 5.07 1.39e-57 6.64e-54 2.66e-13 1.76e-64 4.51e-05 1.99e-04
Llama-3.1-70B 16.44 18.58 16.76 21.40 5.08 5.79 6.50e-42 3.15e-50 3.56e-43 4.20e-61 1.90e-04 5.20e-06
Llama-3.2-1B 0.00 12.97 0.00 11.94 3.25 1.00 1.48e-28 1.00 1.14e-19 1.00
Llama-3.2-11B 15.13 13.25 8.32 12.94 7.41 2.52 7.45e-37 9.27e-30 7.71e-13 1.34e-28 6.34e-10 1.00
Llama-3.2-90B 16.71 17.13 15.74 19.81 5.38 6.32 5.59e-43 1.38e-44 3.23e-39 5.30e-55 4.13e-05 2.63e-07
Qwen-2.5-72B 12.96 17.09 16.79 20.88 -1.12 0.42 1.16e-28 1.90e-44 2.70e-43 3.97e-59 1.00 1.00
Mistral-small 25.53 21.12 27.01 22.72 -0.08 0.02 1.65e-76 4.81e-60 7.12e-82 4.28e-66 1.00 1.00
Mistral-large 24.04 16.25 10.43 23.27 -0.71 1.33 4.76e-71 3.48e-41 1.27e-19 7.47e-68 1.00 1.00
Claude-3 17.07 11.82 17.69 19.29 1.52 -0.44 2.31e-44 1.65e-24 8.75e-47 5.68e-53 1.00 1.00
Claude-3.5-v1 18.09 18.54 22.78 21.61 1.57 2.54 2.47e-48 4.67e-50 2.43e-66 6.55e-62 1.00 1.00
Claude-3.5-v2 25.11 17.45 20.90 20.97 -2.23 0.40 5.34e-75 7.43e-46 3.40e-59 1.93e-59 1.00 1.00
OQA OLMo-2-7B-Base 8.75 11.18 13.23 15.47 -0.46 5.71 5.85e-15 4.04e-24 6.33e-33 1.89e-43 1.00 5.03e-06
OLMo-2-7B-SFT 13.85 11.11 12.47 12.70 15.24 12.01 9.37e-36 7.87e-24 1.54e-29 1.52e-30 2.38e-42 1.51e-27
OLMo-2-7B-DPO 15.60 10.48 7.80 13.52 17.40 12.41 4.52e-44 2.50e-21 7.52e-12 3.16e-34 4.37e-53 3.09e-29
OLMo-2-7B-Instruct 28.64 20.04 15.73 18.59 23.56 14.21 3.07e-113 2.64e-65 9.58e-45 4.99e-55 1.68e-81 5.79e-36
Llama-3-70B 17.00 24.45 21.35 32.48 5.46 6.98 4.83e-51 1.37e-90 7.73e-74 1.32e-133 2.00e-05 2.16e-09
Llama-3.1-70B 36.79 24.62 27.93 13.35 4.98 4.88 9.31e-156 1.63e-91 2.17e-109 2.21e-19 2.42e-04 4.03e-04
Llama-3.2-1B 6.98 20.50 0.00 13.20 0.12 2.17e-09 1.05e-66 1.00 1.52e-27 1.00
Llama-3.2-11B 9.69 14.83 13.04 14.61 9.04 1.26 2.66e-18 2.20e-40 4.33e-32 2.82e-39 3.13e-12 1.00
Llama-3.2-90B 23.61 20.52 21.04 26.79 2.05 4.21 5.28e-86 2.23e-69 3.85e-72 2.91e-103 1.00 8.59e-03
Qwen-2.5-72B 23.40 25.33 34.58 29.76 -0.93 -1.09 7.32e-85 2.52e-95 1.59e-144 3.04e-119 1.00 1.00
Mistral-small 25.22 25.02 24.07 28.47 1.40 1.91 9.29e-95 1.20e-93 1.74e-88 2.50e-112 1.00 1.00
Mistral-large 19.53 23.38 16.79 31.89 1.78 1.19 3.66e-64 9.49e-85 5.54e-50 8.59e-129 1.00 1.00
Claude-3 20.60 20.89 36.32 25.30 -1.69 -1.16 8.62e-70 5.03e-71 2.16e-153 8.09e-95 1.00 1.00
Claude-3.5-v1 22.84 26.23 26.93 32.96 -0.88 -0.48 7.12e-82 3.15e-100 5.45e-104 3.96e-136 1.00 1.00
Claude-3.5-v2 23.75 21.06 23.16 28.23 -0.12 -0.76 8.61e-87 2.77e-72 1.33e-83 5.29e-111 1.00 1.00
WVS OLMo-2-7B-Base 7.51 3.25 10.70 3.59 -5.39 1.06 5.90e-11 3.50e-01 3.11e-22 1.04e-01 2.89e-05 1.00
OLMo-2-7B-SFT 8.88 7.02 9.70 7.42 0.05 3.82 2.05e-15 1.67e-09 2.16e-18 1.12e-10 1.00 4.35e-02
OLMo-2-7B-DPO 14.79 9.23 8.49 10.15 -5.32 2.01 2.75e-40 1.32e-16 4.16e-14 6.37e-20 5.68e-05 1.00
OLMo-2-7B-Instruct 6.14 16.59 5.36 12.43 -0.13 4.43 4.12e-07 1.05e-45 3.42e-05 1.02e-26 1.00 3.39e-03
Llama-3-70B 20.86 14.21 21.30 15.21 2.34 2.89 1.68e-71 2.45e-37 6.87e-74 5.51e-42 1.00 1.00
Llama-3.1-70B 22.28 11.97 17.83 14.92 12.25 6.48 3.50e-79 9.47e-25 1.67e-55 1.55e-35 1.23e-28 1.53e-07
Llama-3.2-1B 0.00 8.56 0.00 5.52 1.00 1.25e-12 1.00 6.97e-05
Llama-3.2-11B 9.13 3.82 9.60 2.80 2.05 2.06 2.59e-16 4.76e-02 5.43e-18 1.00 1.00 1.00
Llama-3.2-90B 16.67 16.70 18.73 18.95 13.14 9.84 1.46e-49 2.28e-49 3.48e-60 7.55e-61 1.44e-32 1.50e-18
Qwen-2.5-72B 38.70 28.16 35.38 27.84 3.54 6.38 1.28e-166 5.87e-110 9.59e-150 3.15e-108 1.27e-01 1.18e-07
Mistral-small 31.25 24.70 28.92 27.31 6.01 9.69 5.40e-128 2.01e-91 1.82e-115 2.05e-105 9.30e-07 5.10e-18
Mistral-large 22.67 19.72 29.96 24.61 6.85 4.80 2.60e-81 8.13e-65 4.52e-121 6.80e-90 5.23e-09 5.94e-04
Claude-3 22.23 13.11 15.77 12.80 5.56 6.99 6.10e-79 3.65e-28 4.98e-45 9.07e-27 1.17e-05 2.47e-09
Claude-3.5-v1 28.00 6.57 25.12 8.89 5.72 2.96 1.90e-110 4.23e-07 9.96e-95 2.37e-12 4.72e-06 1.00
Claude-3.5-v2 26.18 10.74 25.44 13.76 5.56 3.24 1.62e-100 1.64e-20 1.82e-96 3.11e-31 1.15e-05 3.96e-01
Table 14: The t-test statistic and corrected p-value (after Bonferroni correction with alpha 0.000172) between the uncalibrated and calibrated pair for each dataset-model-elicitation method setting. These values are for the results in Appendix Table 7 (subset in the main paper Table 1).
Appendix H Calibration hyperparameters
We perform grid search over the following hyperparameters, tuned on our development set:
  • model: {random forest regression, ridge linear regression, lasso linear regression}
  • random forest
    • num estimators: {100, 125, 150, 175, 200, 225, 250}
    • max depth: {1, 2, 3, 4, 5, 6, None}
  • linear models
    • alpha: {0.1, 0.5, 1, 2, 5, 10, 20, 30, 100, 200}
    • max iterations: {50, 100, 200, 300, 1000, 2000}
Appendix I Prompts
I.1 Demographics used
Demographics are shown in Table 15.
Dataset Qs Example question Demographics
How much do you trust doctors and nurses in your country?
I.2 Base prompts
Base prompts are shown in Table 16.
Prompt type Prompt template
I.3 Sociodemographic prompts
Sociodemographic prompts are shown in Table 17.
Prompt type Prompt template
I.4 Prompt variations
Prompt variations are shown in Table 18.
Prompt variation Prompt template
Table 18: Prompt templates for the different prompt variants. This text is combined with the templates for both base prompts (Table 16) and SD prompts (Table 17) to get the results shown in Tables 4 and 5. All other results use the standard prompts without variation.